Find the value of the Integral
upper limit is pi/2
lower limit is -pi/2
f(x)=root of( cos x - cos^3 x)

Grade:11

1 Answers

kumar
10 years ago
F(x)=root of(cos x -cos^3x) integral of root of(cos x -cos^3x)=(cos x -cos^3x)^3/2 if we put interval then {(cos (pi/2) -cos 3(pi/2)}^3/2- {(cos (-pi/2) -cos 3(-pi/2)}^3/2( ={o-1}^3/2-{o-(-1)}^3/2 =-1-1 =-2 is the answer